OpenStudy (spring98):

A pencil at a stationery store costs $1, and a pen costs $1.50. Stella spent $21 at the store. She bought a total of 18 items. Which system of equations can be used to find the number of pencils (x) and pens (y) she bought? A. x + 18y = 21 x = 1.5y B. 18x + y = 21 x = 1.5y C. x + 1.5y = 21 x + y = 18 D. 1.5x + y = 21 x = 18y

OpenStudy (texaschic101):

she bought a total of 18 items (both pens and pencils) x + y = 18 pencils (x) cost $1 and pens (y) cost $1.50.....totaling $21 1x + 1.50y = 21.....or you could leave the 1 off the x and the 0 off of 1.50 ......x + 1.5y = 21

OpenStudy (spring98):

so is it C?

OpenStudy (spring98):

@texaschic101

OpenStudy (haseeb96):

@Spring98 do you have more questions?

OpenStudy (spring98):

no

OpenStudy (haseeb96):

okay